Solve the system for y:
6x-y = -3
6x-2y = 2

Answer by sarah_adam(201)

6x-y = -3 --- eq 1
6x-2y = 2 --- eq 2
from eq 1: y = 6x+3
from eq 2: 2y = 6x-2 which implies y = 3x-1
now 6x+3 = 3x -1
6x-3x = -3 -1
3x = -4
x = -4/3
but from eq 1
y = 6x+3
y = 6(-(4/3))+3
y = -8+3 = -5
x = -(4/3) and y = -5